cookiemonster:

Which of the following is the most appropriate unit to describe the time you can drive a car based on the amount of gas? Minutes per gallon, because the independent quantity is the gallons Minutes per gallon, because the independent quantity is the minutes Gallons per minute, because the independent quantity is the gallons Gallons per minute, because the independent quantity is the minutes

cookiemonster:

@dude

Vocaloid:

first determine the "independent" and "dependent" variables

Vocaloid:

does the time depend on the number of gallons, or does the number of gallons depend on the time?

cookiemonster:

i think c?

Vocaloid:

think about my question first

cookiemonster:

gallons depends on the time..

Vocaloid:

which do you have direct control over? the number of gallons you put into your car, or the passage of time?

cookiemonster:

number of gallons you put

Vocaloid:

good so the gallons is the independent variable and the time driven is dependent

Vocaloid:

we write rates as "dependent variable per independent variable"

Vocaloid:

so it would be "minutes driven per gallon"

cookiemonster:

ohh b~!

Vocaloid:

almost

Vocaloid:

independent is gallons remember

cookiemonster:

A

Vocaloid:

good

cookiemonster:

:) Which of the following best represents the average rate at which infants grow in height during the first four years after birth? 3 inches per month 3 inches per year 3 feet per month 3 feet per year

Vocaloid:

there are at least 2 that are a bit too fast to be reasonable, which ones do you think they are?

cookiemonster:

i think its b :)

Vocaloid:

good, b is the best answer

cookiemonster:

Which of the following is the most appropriate unit to describe the rate at which people are entering a theme park? Hours per person, because the dependent quantity is the hours Hours per person, because the dependent quantity is the people People per hour, because the dependent quantity is the hours People per hour, because the dependent quantity is the people

Vocaloid:

which is the independent variable? the time Or the number of people?

cookiemonster:

TIME

Vocaloid:

good, so if we write our rate as "dependent variable per independent" which answer matches best?

cookiemonster:

D?

Vocaloid:

good, d is correct

cookiemonster:

A square has a side of 6.25 feet. What is the area in square feet? Round to the correct significant digits. Answer for blank 1:_________

Vocaloid:

area = side^2 = ?

cookiemonster:

39.0625 ??

Vocaloid:

good, but please round that to 3 digits

cookiemonster:

um 39.1?

Vocaloid:

good, that's it

cookiemonster:

Which of the following describes the correct process for solving the equation 2x + 6 = 22 and arrives at the correct solution? Divide both sides by 6 and then add 2. The solution is x = 8. Subtract 6 from both sides and then divide by 2. The solution is x = 8. Add 6 to both sides of the equation and then divide by 2. The solution is x = 6. Subtract 6 from both sides of the equation and then divide by 22. The solution is x = 16.

Vocaloid:

hint: if we have 2x + 6 how can we get 2x by itself and "get rid of" the + 6?

cookiemonster:

b!!!?

Vocaloid:

good, b is correct

cookiemonster:

Name the algebraic property demonstrated in the example below: (2 + x) + y = 2 + (x + y) Commutative Property of Addition Associative Property of Addition Distributive Property Transitive Property

cookiemonster:

A?

Vocaloid:

when you move the parentheses, but not the terms, it's associative

cookiemonster:

oh ok so b?

Vocaloid:

good

cookiemonster:

Name the algebraic property demonstrated in the example below: x ⋅ y ⋅ z = y ⋅ x ⋅ z Distributive Property Transitive Property Associative Property of Multiplication Commutative Property of Multiplication

Vocaloid:

hint: it's the terms that are being moved around, not the parentheses so which one do you think it is?

cookiemonster:

b?

cookiemonster:

no sorry d

Vocaloid:

good, d is correct

cookiemonster:

An equation and the steps to solve it are shown below: 2x − one over four (12x + 4) = 3 Step 1: 2x + 3x + 1 = 3 Step 2: 5x + 1 = 3 Step 3: 5x + 1 − 1 = 3 − 1 Step 4: 5x = 2 In which step was the error made and why? Step 1, because incorrect signs have been used Step 2, because the terms have been added incorrectly Step 3, because a number has been subtracted from both sides of the equation No error was made

cookiemonster:

d?

Vocaloid:

notice how it says 2x − one over four (12x + 4) = 3 with a minus sign in front of the 1/4

Vocaloid:

however, step 2 has a positive sign in front of the 3x + 1

Vocaloid:

meaning they missed a sign in step 1

cookiemonster:

oh ok Which of the following is the most appropriate unit to describe the rate at which data is transferred using the internet? Megabits per second, because the independent quantity is time in seconds and dependent quantity is amount of data in megabits Seconds per megabit, because the independent quantity is time in seconds and dependent quantity is amount of data in megabits Seconds per megabit, because the iIndependent quantity is amount of data in megabits and dependent quantity is time in seconds Megabits per second, because the independent quantity is amount of data in megabits and dependent quantity is time in seconds

Vocaloid:

which one is the independent? megabits or time?

cookiemonster:

megabits

Vocaloid:

in most cases, time is independent (since other variables tend to depend on time)

cookiemonster:

oh ok...

Vocaloid:

(there are some exceptions though)

Vocaloid:

so it should be megabits per second

cookiemonster:

d?

Vocaloid:

time is independent

cookiemonster:

oh sorry A

Vocaloid:

yes good

cookiemonster:

Jon and Lexi are playing a marble game. According to their rules, 4 points are awarded for each red marble collected and 7 points are awarded for each blue marble collected. Which expression represents Jon's total score if he has r red marbles and b blue marbles? 4r − 7b 2r − 7b 11rb 4r + 7b

cookiemonster:

d?

Vocaloid:

yes good

cookiemonster:

Laura is baking a cake. The recipe says that she has to mix 64 grams of chocolate powder to the flour. Laura knows that 1 cup of that particular chocolate powder has a mass of 128 grams. She added two over three of a cup of chocolate powder to the flour. Should Laura add more chocolate powder to make the exact recipe, or did she go over and by what amount? She went over by one over three of a cup She needs to add one over three of a cup She needs to add one over six of a cup She went over by one over six of a cup

Vocaloid:

hint: if 1 whole cup is 128 grams then how many grams is (2/3) of a cup?

cookiemonster:

d!

cookiemonster:

she went over by 1/6

Vocaloid:

yes that's right

cookiemonster:

A truck is rented at $40 per day plus a charge per mile of use. The truck traveled 15 miles in one day, and the total charge was $115. Which equation can be used to find the charge x, in dollars, per mile traveled? 40x + 15 = 115 40 + 15x = 115 40x + 15x = 115 40 + 15 = 115x

Vocaloid:

40 + (number of miles)*(charge per mile) = 115 which equation matches best?

cookiemonster:

c?

Vocaloid:

it's 40 not 40x

cookiemonster:

sorry i mean b

Vocaloid:

yes good

cookiemonster:

Sarah decides to take her six year old son to the circus. The price for the child's ticket is 4.75 dollars less than the price for the adult's ticket. If you represent the price for the child's ticket using the variable "x," how would you write the algebraic expression for the adult's ticket price? 4.75x x + 4.75 x over 4.75 x4.75

Vocaloid:

hint: is the adult's ticket more or less than the child's ticket?

cookiemonster:

more

Vocaloid:

good so does "$4.75 more" imply addition or multiplication?

cookiemonster:

umm addition?

Vocaloid:

good so which answer shows something being added to 4.75?

cookiemonster:

b

Vocaloid:

good

cookiemonster:

An equation is shown below: 4(x − 3) − 5(x + 1) = 3 Which statement shows a correct next step in solving the equation? The equation can become 4x − 3 − 5x + 1 = 3 by applying the associative property of multiplication. The equation can become 4x − 3 − 5x + 1 = 3 by applying the distributive property. The equation can become 4x − 12 − 5x − 5 = 3 by applying the distributive property. The equation can become 4x − 12 − 5x − 5 = 3 by applying the associative property of multiplication.

Vocaloid:

hint: how can we get from − 5(x + 1) to -5x - 5?

Vocaloid:

which property?

cookiemonster:

A

Vocaloid:

try one more time

cookiemonster:

D

Vocaloid:

don't look at the answer choices, just think about which property can get from − 5(x + 1) to -5x - 5?

cookiemonster:

C

Vocaloid:

good, it's the distributive property

cookiemonster:

3 + x = 4 is the same as 4 = 3 + x. This is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

cookiemonster:

C?

Vocaloid:

notice how the positions of the terms aren't changing, the entire arrangement is being "reflected" A = B then B = A is the symmetric property

Vocaloid:

like a mirror

cookiemonster:

oh ok

cookiemonster:

Ruth borrowed a book from a library. The library charged a fixed rental for the book and a late fee for every day the book was overdue. The expression below shows the charges Ruth paid for the book when she returned it x days after the due date: 2 + 0.25x What does the second term of the expression represent? The total late fee paid The fixed rental for the book The late fee per day for the book The number of days the book was overdue

Vocaloid:

since 0.25x has a variable, it must involve the number of days somehow

Vocaloid:

0.25 is the late fee (per day) x is the number of days so what do you get when you multiply these together?

cookiemonster:

c?

Vocaloid:

we said that 0.25 is the late fee per day

Vocaloid:

but what does 0.25x represent?

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid 0.25 is the late fee (per day) x is the number of days so what do you get when you multiply these together? \(\color{#0cbb34}{\text{End of Quote}}\)

cookiemonster:

ummm...

Vocaloid:

if the late fee per day is $0.25 and the number of days is x how would you calculate the total late fee paid?

cookiemonster:

multiply?

Vocaloid:

good, so 0.25x = total late fee paid

Vocaloid:

which answer matches this description?

cookiemonster:

A

Vocaloid:

good

cookiemonster:

Kevin plans to put concrete on a rectangular portion of his driveway. The portion is 10 feet long and 3 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Kevin needs is $72.59. Which of the following is closest to the width of the portion of the driveway on which Kevin plans to put concrete? [1 foot = 12 inches; 1 yard = 3 feet] 3 feet 6 feet 8 feet 10 feet

Vocaloid:

if the total cost is $72.59 and each cubic yard of concrete costs $98 how many cubic yards of concrete did he buy?

cookiemonster:

6

Vocaloid:

how did you get that answer?

cookiemonster:

no like the answer is B :)

Vocaloid:

again, how did you get that answer?

cookiemonster:

omg! no wait i did it wrong woops you have to divide.. $72.59 / $98.. then you get 0.74.. This is equal to 19.98 cubic feet. This means the width of the portion is equal to 8 feet.

Vocaloid:

good, 8 is your answer

cookiemonster:

Which phrase matches the algebraic expression below? 3(x − 7) + 10 Three times the sum of x and seven plus ten Three times x minus seven plus ten Three times x minus the sum of seven and ten Three times the difference of x and seven plus ten

cookiemonster:

btw, random question... after im done with my math questions can you help me with a history assignment about the lewis and clark expedition?

Vocaloid:

hint: how would you write "(x − 7)" in words?

cookiemonster:

x minus seven?

Vocaloid:

good, or "the difference of x and 7"

Vocaloid:

now, 3(x-7) is just 3 times that difference

Vocaloid:

so the entire expression is Three times the difference of x and seven plus ten

cookiemonster:

so d?

Vocaloid:

yes

cookiemonster:

ok done with the math questions